OpenStudy (anonymous):

Given the following endpoints (A & B) and midpoint (M) on a number line, find the coordinates for A & B. M(-11) A(2x + 3) B(5x + 3) 1) A(-13) & B(-9) 2) A(-5) & B(-17) 3) A(-17) & B(-5) 4) A(-9) & B(-13)

OpenStudy (anonymous):

it exists (A+B)/2=M so (2x+3+5x+3)/2=-11 3.5x+3=-11, 3.5x=-14, x=-4 so A=-5 B=-17

OpenStudy (anonymous):

Thanks AndrewNJ